On December 31, 2017, Berclair Inc. had 600 million shares of common stock and 7 million shares of 9%, $100 par value cumulative preferred stock issued and outstanding. On March 1, 2018, Berclair purchased 184 million shares of its common stock as treasury stock. Berclair issued a 5% common stock dividend on July 1, 2018. Four million treasury shares were sold on October 1. Net income for the year ended December 31, 2018, was $1,150 million.

Also outstanding at December 31 were 30 million incentive stock options granted to key executives on September 13, 2013. The options were exercisable as of September 13, 2017, for 30 million common shares at an exercise price of $56 per share. During 2018, the market price of the common shares averaged $70 per share.

Required:

Compute Berclair's basic and diluted earnings per share for the year ended December 31, 2018
Answer #1
Berclair inc.
Basic EPS = ( net incoome - preference dividend ) / Weighted average number of common shares outstanding.
Weighted average number of common shares outstanding = 600 * 12/12 * 1.05 - 184 * 10/12 * 1.05 + 4 * 3/12 = 630 - 161 + 1 = 470 million shares
( 1150 - 7 * 100 * 9 % ) / 470 = 1087 / 470 = $ 2.313 per share
Dilutive EPS
Exercisable securities :- Use of treasury stock if dilutive
The securities are dilutive since the exercise price of $ 56 is less than the average market price of $70.
Assume exercise date to be 1/1/17
Proceeds recived on exercise = 30 * 56 = 1680 million
Shares repurchased = 1680 / 70 = 24 million
Net increase in shares outstanding = 30 - 24 = 6 million
EPS with inclusion of options :- 1087 / 470 + 6 = $ 2.284
Required answer
Numerator / Denominator = EPS
Basic 1087/ 470 = 2.313
Diluted 1087/ 476 = 2.284